# Kinetic-Ballooning-Bifurcation in Tokamak Pedestals Across Shaping and Aspect-Ratio

### 08 Dec 2023 v1

@@ Abstract
Branch-Jumping. – Finally, we note a curious consequence of the width-height bifurcation: a sufficiently big and sudden loss (or gain) of pedestal particles and heat – such as an ELM (or pellet injection) – could cause a pedestal to jump from the narrow to wide branch (or wide to narrow branch) in Figure 6. There are examples of H-modes that become ELM-free after an initial ELM [37]; jumping from narrow to wide branches is a possible explanation for these observations. Finally, two companion NSTX discharges shown in Figure 1, 129015 and 129038 [14], fall in the the narrow and wide branches. These discharges differ mainly by the amount of Lithium dosing, which might ultimately be responsible for different branch access.
@@
Summary. – We have discovered a novel and intriguing bifurcation in the width-height scalings of tokamak pedestals, which can be significantly modified by plasma shaping and aspect-ratio. This bifurcation arises from the interplay between first and second stability properties of kinetic-ballooning-modes and allows for the existence of both wide and narrow pedestal branches. Importantly, this discovery opens up the operating space for accessible pedestal widths and heights, offering new prospects for pedestal control in current and future fusion experiments. While our investigations thus far have only revealed wide branch access in NSTX pedestals, investigation of how NSTX pedestals accessed the wide branch could provide a pathway to ELM-free operation while achieving high pedestal pressures.
@@
Acknowledgements. – We are grateful to S. C. Cowley, W. Dorland, R. Maingi, O. Sauter, P. B. Snyder, and H. R. Wilson for insightful discussions. This work was supported by the U.S. Department of Energy under contract numbers DE-AC02-09CH11466, DE-SC0022270, DE-SC0022272, and the Department of Energy Early Career Research Program. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.

### 07 Apr 2024 v2

@@ Abstract
While we lack a prescription for accessing separate branches of the width-height scaling bifurcation, there is a curious observation: a sufficiently big and sudden loss (or gain) of pedestal particles and heat – such as an ELM (or pellet injection) – could cause a pedestal to jump from the narrow to wide branch (or wide to narrow branch) in fig. 7. There are examples of H-modes that become ELM-free after an initial ELM [52]; jumping from narrow to wide branches is a possible explanation. Finally, companion NSTX discharges 129015 and 129038 [23], shown in fig. 1, fall in the the narrow and wide branches. These discharges differ mainly by Lithium dosing [53], which may be responsible for different branch access. Two other largely ELM-free regimes, I-mode [54] and wide pedestal QH-mode [55], may also access the wide pedestal branch.
@@
We have discovered a novel and intriguing bifurcation in the width-height scalings of tokamak pedestals, which can be modified by plasma shaping and aspect-ratio [33]. This bifurcation arises from the first and second stability properties of kinetic-ballooning-modes that give distinct wide and narrow pedestal branches. This discovery opens up the operating space for accessible pedestal widths and heights, offering new prospects for pedestal regimes in fusion experiments. While our investigations thus far have only revealed wide branch access in NSTX pedestals, investigation of how NSTX pedestals access the wide branch could provide a pathway to ELM-free operation [3] with high pedestal pressures.
@@
We are grateful to S. C. Cowley, W. Dorland, R. Maingi, O. Sauter, P. B. Snyder, and H. R. Wilson for insightful discussions and to T. Bechtel and J. McClenaghan for technical assistance with the EFUND code. This work was supported by the U.S. Department of Energy under contract numbers DE-AC02-09CH11466, DE-SC0022270, DE-SC0022272, DE-SC0014264, DE-SC0021629, and the Department of Energy Early Career Research Program. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
The data that support the findings of this study are openly available in Princeton Data Commons at https://doi.org/10.34770/fc813051. Part of the data analysis was performed using the OMFIT integrated modeling framework [56] using the Github project gk_ped [15].